# Fourier interpolation in dimensions 3 and 4 and real-variable Kloosterman sums

### v1

@@ 1. Introduction
    $|B_{4,n}(r)|,|\widetilde{B}_{4,n}(r)|\ll_{\varepsilon}\left\{\begin{array}[]{ll}n^{1+\varepsilon}&\text{if }r^{2}\leq 1/n,\\ \max(|r|^{-1},|r|^{-\frac{1}{4}})n^{\frac{3}{4}+\varepsilon}&\text{if }r^{2}\geq 1/n.\end{array}\right$.
@@
    The values $B_{4,n}(0)$ and $\widetilde{B}_{4,n}(0)$ are given by
@@
    Here $\kappa\in[0,1]$ is chosen so that $2^{\nu_{2}(n)}\ll n^{\kappa}$, where $\nu_{2}(n)$ is the 2-adic valuation of $n$. Note that $\frac{437}{588}\approx 0.743\ldots$.
@@ 3.4. Computation of Fourier expansions
$\begin{array}[]{r}|B_{4,n}(r;s,y)|\vskip 5.01874pt\\ |\widetilde{B}_{4,n}(r;s,y)|\end{array}\ll_{\varepsilon}\left\{\begin{array}[]{ll}n^{1+\varepsilon}e^{-\pi yn}&\text{if }n\geq 1,\ r^{2}n\leq 1,\\ (|r|^{-1}+|r|^{-\frac{1}{2}})n^{\frac{3}{4}+\varepsilon}e^{-\pi yn}&\text{if }n\geq 1,\ r^{2}n\geq 1,\\ (|r|^{-1}+|r|)|n|e^{-\pi y|n|+2\pi|r||n|^{1/2}}/\Gamma(s-1)&\text{if }n\leq-1.\end{array}\right$.
@@ 4. Kloosterman sums in weight 2
Moreover, by [ST09, Theorem 2, Corollary 3], for $mn\neq 0$ we have
$\sum_{1\leq c\leq x}\frac{S(m,n,c)}{c}\ll_{\varepsilon}\left(x^{\frac{1}{6}}+|mn|^{\frac{1}{6}}+(|m|+|n|)^{\frac{1}{8}}|mn|^{\frac{7}{128}}\right)|xmn|^{\varepsilon}.$ (4.2)
In the proofs we will need to deal with the sum above on $4|c\leq x$ and $8|c\leq x$. Such an estimate for a sum of Kloosterman sums over an arithmetic progression $c\equiv 0\ (\mathrm{mod}\ N)$, $c\leq x$ has been proved by [GS11, Theorem 2, Corollary 3] with an extra term added in the erratum [GS20, Theorem 1, Corollary 1]. Note that the extra term in the erratum is due to exceptional eigenvalues of Laplacian on $\Gamma_{0}(N)$ in general. By Lemma 2.7, there is no exceptional eigenvalue when $N=4,8$ and the bound in [GS11, Corollary 3] remains valid for our application. Moreover, for the case $mn<0$, the bound [GS11, Corollary 3] also remains valid, since in the trace formula we do not have the case $m=n$ (see [GS11, Theorem 5]).
Lemma 4.1.
Suppose $mn\neq 0$. For $N=4$ or $N=8$, we have
$\sum_{N|c\leq x}\frac{S(m,n,c)}{c}\ll_{\varepsilon}\left(x^{\frac{1}{6}}+|mn|^{\frac{1}{6}}+(|m|+|n|)^{\frac{1}{8}}|mn|^{\frac{7}{128}}\right)|xmn|^{\varepsilon}.$ (4.3)
By applying (2.14) and (2.20), the following estimate follows from Lemma 4.1.
Lemma 4.2.
For $mn\neq 0$, we have
$\left.\begin{array}[]{r}\displaystyle\sum_{2|c\leq x}\frac{S(m,n,c,\nu_{\Theta}^{4})}{c}\\ \displaystyle\sum_{2\nmid\widetilde{c}\leq x}\frac{S(m,n,\widetilde{c},\nu_{\Theta}^{4})}{\widetilde{c}}\end{array}\right\}\ll_{\varepsilon}\left(x^{\frac{1}{6}}+|mn|^{\frac{1}{6}}+(|m|+|n|)^{\frac{1}{8}}|mn|^{\frac{7}{128}}\right)|xmn|^{\varepsilon}.$ (4.4)
@@ 5.1. Case $n\neq 0$
For $k\in{\mathbb{Z}}$, the Weil-type bound is by the Weil bound of standard Kloosterman sums and (2.14):
$|S(k,n,c,\nu_{\Theta}^{4})|=|S(k,n,2c)|\leq\sigma_{0}(2c)\sqrt{(k,n,2c)}\sqrt{2c}.$ (5.1)
Then by Lemma 2.8, we have
@@
For $x\geq 1,r\in{\mathbb{R}}_{+}$ and $n\neq 0$, let $X=\max(x,r)$. We have the following estimates:
$\left.\begin{array}[]{r}\displaystyle\sum_{2|c\leq x}\frac{K(r,n,c,\nu_{\Theta}^{4})}{c}\\ \displaystyle\sum_{2\nmid\widetilde{c}\leq x}\frac{K(r,n,\widetilde{c},\nu_{\Theta}^{4})}{\widetilde{c}}\\ \end{array}\right\}\ll_{\varepsilon}|Xn|^{\frac{23}{128}+\varepsilon}.$ (5.3)
Remark.
In the proof we can get a better uniform bound
$\ll_{\varepsilon}\left((X+|n|)^{\frac{1}{8}}|Xn|^{\frac{7}{128}}+|Xn|^{\frac{1}{6}}\right)|Xn|^{\varepsilon},$
but our goal is to focus on the growth rate when $|n|$ is large. This is the reason for a slightly crude bound, which simplifies the further applications.
@@
The case $r\in{\mathbb{Z}}$ follows from the better bounds in §4, so we only prove the case $r\notin{\mathbb{Z}}$ here.
Let $K_{4e}(x)$ denote the first line of the left hand side of (5.3). We only prove the bound for $K_{4e}(x)$ and the other case is similar.
We apply Lemma 2.8 to get
$K_{4e}(x)=\sum_{2|c\leq x}\sum_{k(2c)}\frac{i\sin(\pi r)(-1)^{k}}{{\mathbf{e}}(\frac{r-k}{2c})-1}\cdot\frac{S(k,n,c,\nu_{\Theta}^{4})}{c^{2}}.$
If we define
$S_{4e}(x)\vcentcolon=\sum_{2|c\leq x}\frac{S(k,n,c,\nu_{\Theta}^{4})}{c^{2}},$
then by (4.4) and partial summation, for $k\neq 0$ we have
$S_{4e}(x)\ll_{\varepsilon}\left(x^{-\frac{5}{6}}+\frac{|kn|^{\frac{1}{6}}}{x}+\frac{(k+|n|)^{\frac{1}{8}}|kn|^{\frac{7}{128}}}{x}\right)|xkn|^{\varepsilon}.$ (5.4)
Then,
$\begin{split}\frac{K_{4e}(x)}{i\sin(\pi r)}=&\sum_{2|c\leq x}\sum_{\begin{subarray}{c}k\in{\mathbb{Z}}\\ |r-k|<c\end{subarray}}(-1)^{k}\frac{S(k,n,2c)}{{\mathbf{e}}(\frac{r-k}{2c}-1)c^{2}}\\ =&\sum_{\begin{subarray}{c}2|c\\ r<c\leq x\end{subarray}}\frac{S(0,n,2c)}{\left({\mathbf{e}}(\frac{r}{2c})-1\right)c^{2}}\\ &+\sum_{\begin{subarray}{c}0\neq k\in{\mathbb{Z}}\\ r-x\leq k<r\\ \end{subarray}}(-1)^{k}\sum_{\begin{subarray}{c}2|c\\ r-k<c\leq x\end{subarray}}\frac{S(k,n,2c)}{\left({\mathbf{e}}(\frac{r-k}{2c})-1\right)c^{2}}\\ &+\sum_{\begin{subarray}{c}k\in{\mathbb{Z}}\\ r<k\leq r+x\end{subarray}}(-1)^{k}\sum_{\begin{subarray}{c}2|c\\ k-r<c\leq x\end{subarray}}\frac{S(k,n,2c)}{\left({\mathbf{e}}(\frac{r-k}{2c})-1\right)c^{2}}\\ =\vcentcolon&\Sigma_{4e0}+\Sigma_{4e1}+\Sigma_{4e2}.\end{split}$ (5.5)
Here $\Sigma_{4e0}$ appears only when $x>r$. The estimates for $\Sigma_{4e0}$, $\Sigma_{4e1}$ and $\Sigma_{4e2}$ are similar. We begin with $\Sigma_{4e2}$ where $r<k\leq r+x$. By (5.2), we have
$\sum_{\begin{subarray}{c}2|c\\ k-r<c\leq x\end{subarray}}$ $\frac{S(k,n,2c)}{({\mathbf{e}}(\frac{r-k}{2c})-1)c^{2}}=\int_{k-r+\|r\|/2}^{x}\frac{dS_{4e}(t)}{{\mathbf{e}}(\frac{r-k}{2t})-1}$
$\ll\frac{|S_{4e}(x)|}{|{\mathbf{e}}(\frac{r-k}{2x})-1|}+{|S_{4e}(k-r+\tfrac{\|r\|}{2})|}+\left|\int_{k-r+\|r\|/2}^{x}\frac{S_{4e}(t)(r-k){\mathbf{e}}(\frac{r-k}{2t})}{t^{2}\left({\mathbf{e}}(\frac{r-k}{2t})-1\right)^{2}}dt\right|$
$\ll\left(\frac{x^{\frac{1}{6}}}{k-r}+\frac{|kn|^{\frac{1}{6}}}{k-r}+\frac{(k+|n|)^{\frac{1}{8}}|kn|^{\frac{7}{128}}}{k-r}\right)|xkn|^{\varepsilon}$
$+\int_{k}^{x}\frac{t^{-\frac{5}{6}}+t^{-1}|kn|^{\frac{1}{6}}+t^{-1}(k+|n|)^{\frac{1}{8}}|kn|^{\frac{7}{128}}}{k-r}|tkn|^{\varepsilon}dt$
$\ll\left(\frac{x^{\frac{1}{6}}}{k-r}+\frac{|kn|^{\frac{1}{6}}}{k-r}+\frac{(k+|n|)^{\frac{1}{8}}|kn|^{\frac{7}{128}}}{k-r}\right)|xkn|^{\varepsilon}.$
Let $X=\max(x,r)$. Summing up the above on $k$ for $r<k\leq r+x$ and applying Lemma 2.10 we get
$\Sigma_{4e2}\ll_{\varepsilon}\left(|Xn|^{\frac{1}{8}+\frac{7}{128}}+|Xn|^{\frac{1}{6}}\right)\frac{|Xn|^{\varepsilon}}{\|r\|}.$ (5.6)
Similarly, for $\Sigma_{4e1}$ we get
$\sum_{\begin{subarray}{c}k\leq 2c\leq 2x\\ 2|c\end{subarray}}\frac{S(k,n,2c)}{{\mathbf{e}}(\frac{r-k}{2c}-1)c^{2}}\ll\left(\frac{x^{\frac{1}{6}}}{r-k}+\frac{|kn|^{\frac{1}{6}}}{r-k}+\frac{(k+|n|)^{\frac{1}{8}}|kn|^{\frac{7}{128}}}{r-k}\right)|xkn|^{\varepsilon}.$
Summing up on $r-x\leq k<r$ we get the same bound as (5.6). For $\Sigma_{4e0}$ the process is similar by applying (4.5) and we have smaller bound.
Combining (5.5), (5.6), and noticing that $\frac{|\sin(\pi r)|}{\|r\|}\in[2,\pi]$ for $r\in{\mathbb{R}}$, we have proved the bound for $K_{4e}$ in (5.3). ∎
Proposition 5.4.
Let $s\in[1,1.001]$ and $r\in{\mathbb{R}}_{+}$. If $n>0$, both the following sums are convergent and have the same estimate depending on $r^{2}n\leq 1$ or $r^{2}n\geq 1$:
$\text{both}\quad\left.\begin{array}[]{r}\displaystyle\sum_{2|c>0}\frac{K(r^{2},n,c,\nu_{\Theta}^{4})}{c}J_{2s-1}\left(\frac{2\pi|r|\sqrt{n}}{c}\right)\\ \displaystyle\sum_{2\nmid\widetilde{c}>0}\frac{\widetilde{K}(r^{2},n,\widetilde{c},\nu_{\Theta}^{4})}{\widetilde{c}}J_{2s-1}\left(\frac{2\pi|r|\sqrt{n}}{\widetilde{c}}\right)\\ \end{array}\right\}\ll_{\varepsilon}\left\{\begin{array}[]{ll}|r|n^{\frac{1}{2}+\varepsilon},&\text{if }r^{2}\leq\frac{1}{n},\vskip 5.01874pt\\ (1+|r|^{\frac{1}{2}})n^{\frac{1}{4}+\varepsilon},&\text{if }r^{2}\geq\frac{1}{n}.\end{array}\right$.
If $n<0$, the following sums are convergent and have estimates
$\left.\begin{array}[]{r}\displaystyle\sum_{2|c>2\pi|r^{2}n|^{1/2}}\frac{K(r^{2},n,c,\nu_{\Theta}^{4})}{c}I_{2s-1}\left(\frac{2\pi|r^{2}n|^{\frac{1}{2}}}{c}\right)\\ \displaystyle\sum_{2\nmid\widetilde{c}>2\pi|r^{2}n|^{1/2}}\frac{\widetilde{K}(r^{2},n,\widetilde{c},\nu_{\Theta}^{4})}{\widetilde{c}}I_{2s-1}\left(\frac{2\pi|r^{2}n|^{\frac{1}{2}}}{\widetilde{c}}\right)\\ \end{array}\right\}\ll_{\varepsilon}(1+|r|^{\frac{1}{2}})|n|^{\frac{69}{256}+\varepsilon}.$
$\left.\begin{array}[]{r}\displaystyle\sum_{2|c\leq 2\pi|r^{2}n|^{1/2}}\frac{K(r^{2},n,c,\nu_{\Theta}^{4})}{c}I_{2s-1}\left(\frac{2\pi|r^{2}n|^{\frac{1}{2}}}{c}\right)\\ \displaystyle\sum_{2\nmid\widetilde{c}\leq 2\pi|r^{2}n|^{1/2}}\frac{\widetilde{K}(r^{2},n,\widetilde{c},\nu_{\Theta}^{4})}{\widetilde{c}}I_{2s-1}\left(\frac{2\pi|r^{2}n|^{\frac{1}{2}}}{\widetilde{c}}\right)\\ \end{array}\right\}\ll_{\varepsilon}(1+|r|^{2})|n|e^{2\pi|r||n|^{1/2}}.$
Remark.
There are better uniform estimates when $r^{2}$ is large, while we focus more on the $|n|$ aspect with the same reason as the remark under Proposition 5.3.
Proof.
The proof requires Proposition 5.3 and the following bounds on $J$- and $I$-Bessel functions. By [DLMF, (10.14.1)], $|J_{\nu}(x)|\leq 1$ for $\nu\geq 0$; by [DLMF, (10.7.3), (10.7.8)],
$J_{\nu}(x)\ll\nu\min\left(x^{-\frac{1}{2}},x^{\nu}\right),\quad\text{hence }J_{2s-1}(x)\ll\min\left(x^{-\frac{1}{2}},x\right)\text{ for }s\in[1,1.001],$ (5.7)
and by [DLMF, (10.6.1)],
$\begin{split}\frac{d}{dt}J_{2s-1}\left(\frac{2\pi|r|\sqrt{n}}{t}\right)&=\frac{\pi|r|\sqrt{n}}{t^{2}}\left(J_{2s}\left(\frac{2\pi|r^{2}n|^{\frac{1}{2}}}{t}\right)-J_{2s-2}\left(\frac{2\pi|r^{2}n|^{\frac{1}{2}}}{t}\right)\right)\\ &\ll\frac{|r|\sqrt{n}}{t^{2}}.\end{split}$ (5.8)
Let $\beta>\frac{1}{2}$, the precise value of which will be chosen later. By Lemma 5.1 and (5.7), we have the following two cases:
(1) when $r^{2}\leq\frac{1}{n}$, the following sum converges and has estimate
@@
(2) when $r^{2}\geq\frac{1}{n}$, we break the sum at $c\asymp|r^{2}n|^{\beta}$ and get
$\begin{split}&\sum_{2|c,\,c\leq|r^{2}n|^{\beta}}\frac{K(r^{2},n,c,\nu_{\Theta}^{4})}{c}J_{2s-1}\left(\frac{2\pi|r|\sqrt{n}}{c}\right)\\ &\ll_{\varepsilon}|r^{2}n|^{-\frac{1}{4}}\sum_{2|c,\,c\leq|r^{2}n|^{1/2}}(n,c)^{\frac{1}{2}+\varepsilon}c^{\varepsilon}+|r^{2}n|^{\frac{1}{2}}\sum_{\begin{subarray}{c}2|c\\ |r^{2}n|^{1/2}\leq c\leq|r^{2}n|^{\beta}\end{subarray}}(n,c)^{\frac{1}{2}+\varepsilon}c^{-\frac{3}{2}+\varepsilon}\\ &\ll_{\varepsilon}|r^{2}n|^{\frac{1}{4}+\varepsilon}.\end{split}$ (5.10)
By Proposition 5.3, (5.8) and partial summation, for any $y>x>\max(1,r^{2},2\pi|r|\sqrt{n})$,
$\sum_{2|c,\ x\leq c\leq y}\frac{K(r^{2},n,c,\nu_{\Theta}^{4})}{c}J_{2s-1}\left(\frac{2\pi|r|\sqrt{n}}{c}\right)\ll_{\varepsilon}|r|n^{\frac{1}{2}}x^{-\frac{105}{128}+\varepsilon},$ (5.11)
hence the following tail sum is convergent by Cauchy’s criterion and we have
$\sum_{2|c\geq|r^{2}n|^{\beta}}\frac{K(r^{2},n,c,\nu_{\Theta}^{4})}{c}J_{2s-1}\left(\frac{2\pi|r|\sqrt{n}}{c}\right)\ll_{\varepsilon}(|r|^{\beta}+|r|)^{\frac{23}{64}}n^{\frac{23}{128}(1+\beta)+\frac{1}{2}-\beta+\varepsilon}.$ (5.12)
Whenever $\beta\geq\frac{11}{21}$, the exponent of $n$ in (5.12) is dominated by it in (5.10). The proof of the $n>0$ case in Proposition 5.4 follows from combining (5.9) and (5.10).
For the $n<0$ case the proof is also done via partial summation and the following bounds: by [DLMF, (10.30.1)], for fixed $\nu\geq 0$, $I_{\nu}(x)\ll_{\nu}x^{\nu}$ for $x\leq 1$, hence $I_{2s-1}(x)\ll x$ for $x\leq 1$; by [DLMF, (10.30.4)], $I_{\nu}(x)\ll e^{x}$ for $x\geq 1$; by [DLMF, (10.29.1)]
[…]
The convergence follows from Cauchy’s criterion and the bound is derived by combining Proposition 5.3 and the above inequalities.
The proof for case $2\nmid\widetilde{c}$ follows by the same argument. ∎

### v2

@@ 1. Introduction
    $|B_{4,n}(r)|,|\widetilde{B}_{4,n}(r)|\ll_{\varepsilon}\left\{\begin{array}[]{ll}n^{1+\varepsilon}&\text{if }r^{2}\leq 1/n,\\ (|r|^{-1}+|r|^{-\frac{1}{2}+\varepsilon})n^{\frac{3}{4}+\varepsilon}&\text{if }r^{2}\geq 1/n.\end{array}\right$.
@@
    Let $\nu_{2}(n)$ be the 2-adic valuation of $n$. The values $B_{4,n}(0)$ and $\widetilde{B}_{4,n}(0)$ are given by
@@
    Here $\kappa\in[0,1]$ is chosen so that $2^{\nu_{2}(n)}\ll n^{\kappa}$. Note that $\frac{437}{588}\approx 0.743\ldots<0.75$.
@@ Acknowledgement
Conflict of interest statement
On behalf of all authors, the corresponding author states that there is no conflict of interest.
Data availability statement
No data were generated or analyzed in this study.
@@ 3.4. Computation of Fourier expansions
$\begin{array}[]{r}|B_{4,n}(r;s,y)|\vskip 5.01874pt\\ |\widetilde{B}_{4,n}(r;s,y)|\end{array}\ll_{\varepsilon}\left\{\begin{array}[]{ll}n^{1+\varepsilon}e^{-\pi yn}&\text{if }n\geq 1,\ r^{2}n\leq 1,\\ (|r|^{-1}+|r|^{-\frac{1}{2}+\varepsilon})n^{\frac{3}{4}+\varepsilon}e^{-\pi yn}&\text{if }n\geq 1,\ r^{2}n\geq 1,\\ (|r|^{-1}+|r|^{1+\varepsilon})|n|^{2}e^{-\pi y|n|+2\pi|r||n|^{1/2}}/\Gamma(s-1)&\text{if }n\leq-1.\end{array}\right$.
@@ 4. Kloosterman sums in weight 2
By Proposition 2.2, for positive integers $2|c$ and $2\nmid\widetilde{c}$ we have
[…]
@@ 5.1. Case $n\neq 0$
For $k\in{\mathbb{Z}}$, recall the Weil-type bound of Kloosterman sums (4.2) and $\sigma_{0}(n)\ll_{\varepsilon}n^{\varepsilon}$. Then by Lemma 2.8,
@@
Let $s\in[1,1.001]$ and $r\in{\mathbb{R}}_{+}$. If $n>0$, both the following sums are absolute convergent and have the same estimate depending on $r^{2}n\leq 1$ or $r^{2}n\geq 1$:
$\text{both}\quad\left.\begin{array}[]{r}\displaystyle\sum_{2|c>0}\frac{K(r^{2},n,c,\nu_{\Theta}^{4})}{c}J_{2s-1}\left(\frac{2\pi|r|\sqrt{n}}{c}\right)\\ \displaystyle\sum_{2\nmid\widetilde{c}>0}\frac{\widetilde{K}(r^{2},n,\widetilde{c},\nu_{\Theta}^{4})}{\widetilde{c}}J_{2s-1}\left(\frac{2\pi|r|\sqrt{n}}{\widetilde{c}}\right)\\ \end{array}\right\}\ll_{\varepsilon}\left\{\begin{array}[]{ll}|r|n^{\frac{1}{2}+\varepsilon},&\text{if }r^{2}\leq\frac{1}{n},\vskip 5.01874pt\\ (1+|r|^{\frac{1}{2}+\varepsilon})n^{\frac{1}{4}+\varepsilon},&\text{if }r^{2}\geq\frac{1}{n}.\end{array}\right$.
If $n<0$, the following sums are absolute convergent and have estimates
$\left.\begin{array}[]{r}\displaystyle\sum_{2|c>2\pi|r^{2}n|^{1/2}}\frac{K(r^{2},n,c,\nu_{\Theta}^{4})}{c}I_{2s-1}\left(\frac{2\pi|r^{2}n|^{\frac{1}{2}}}{c}\right)\\ \displaystyle\sum_{2\nmid\widetilde{c}>2\pi|r^{2}n|^{1/2}}\frac{\widetilde{K}(r^{2},n,\widetilde{c},\nu_{\Theta}^{4})}{\widetilde{c}}I_{2s-1}\left(\frac{2\pi|r^{2}n|^{\frac{1}{2}}}{\widetilde{c}}\right)\\ \end{array}\right\}\ll_{\varepsilon}(1+|r|^{1+\varepsilon})|n|^{\frac{1}{2}+\varepsilon}.$
$\left.\begin{array}[]{r}\displaystyle\sum_{2|c\leq 2\pi|r^{2}n|^{1/2}}\frac{K(r^{2},n,c,\nu_{\Theta}^{4})}{c}I_{2s-1}\left(\frac{2\pi|r^{2}n|^{\frac{1}{2}}}{c}\right)\\ \displaystyle\sum_{2\nmid\widetilde{c}\leq 2\pi|r^{2}n|^{1/2}}\frac{\widetilde{K}(r^{2},n,\widetilde{c},\nu_{\Theta}^{4})}{\widetilde{c}}I_{2s-1}\left(\frac{2\pi|r^{2}n|^{\frac{1}{2}}}{\widetilde{c}}\right)\\ \end{array}\right\}\ll_{\varepsilon}(1+|r|^{2+\varepsilon})|n|^{1+\varepsilon}e^{2\pi|r||n|^{1/2}}.$
@@
The proof requires the following bounds on $J$- and $I$-Bessel functions. By [DLMF, (10.14.1)], $|J_{\nu}(x)|\leq 1$ for $\nu\geq 0$; by [DLMF, (10.7.3), (10.7.8)],
$J_{\nu}(x)\ll_{\nu}\min\left(x^{-\frac{1}{2}},x^{\nu}\right),\quad\text{hence }J_{2s-1}(x)\ll\min\left(x^{-\frac{1}{2}},x\right)\text{ for }s\in[1,1.001].$ (5.2)
By Lemma 5.1 and (5.2), we have the following two cases:
(1) when $r^{2}\leq\frac{1}{n}$, the following sum converges absolutely and has estimate
@@
(2) when $r^{2}\geq\frac{1}{n}$, we break the sum at $|r^{2}n|^{1/2}$ and still get the absolute convergence
$\begin{split}&\sum_{2|c>0}\frac{K(r^{2},n,c,\nu_{\Theta}^{4})}{c}J_{2s-1}\left(\frac{2\pi|r|\sqrt{n}}{c}\right)\\ &\ll_{\varepsilon}|r^{2}n|^{-\frac{1}{4}}\sum_{2|c,\,c\leq|r^{2}n|^{1/2}}(n,c)^{\frac{1}{2}+\varepsilon}c^{\varepsilon}+|r^{2}n|^{\frac{1}{2}}\sum_{2|c,\,c\geq|r^{2}n|^{1/2}}(n,c)^{\frac{1}{2}+\varepsilon}c^{-\frac{3}{2}+\varepsilon}\\ &\ll_{\varepsilon}|r^{2}n|^{\frac{1}{4}+\varepsilon}.\end{split}$ (5.4)
The case $2\nmid\widetilde{c}$ follows similarly.
For $n<0$ we need following bounds for the $I$-Bessel function: by [DLMF, (10.30.1)], for fixed $\nu\geq 0$, $I_{\nu}(x)\ll_{\nu}x^{\nu}$ for $x\leq 1$, hence $I_{2s-1}(x)\ll x$ for $x\leq 1$; by [DLMF, (10.30.4)], $I_{\nu}(x)\ll e^{x}$ for $x\geq 1$.
Those series are absolute convergent following from Lemma 5.1, Proposition 5.2 and the above inequalities. The proof for case $2\nmid\widetilde{c}$ follows by the same argument. We have finished the proof of the proposition. ∎